In one instance a financial institution loaned you $70,000 for two years at an APR of 8.75%. In a second instance you loaned a financial institution $70,000 for two years at an APR of 8.75% compounded monthly. What is the difference in the amount of interest paid? (Round your answer to the nearest cent.)
I cannot seem to get the correct answer. I think I am doing it incorrectly. Can someone help.

Answer by stanbon(75887)

In one instance a financial institution loaned you $70,000 for two years at an APR of 8.75%.
I = P*r*t = 70,000*0.0875*2 = $12,250
----------------------------
In a second instance you loaned a financial institution $70,000 for two years at an APR of 8.75% compounded monthly.
I = 70,000(1+(.0875/12))^24 - 70,000 = $83,334.31-70,000 = 13,334.31
What is the difference in the amount of interest paid? (Round your answer to the nearest cent.)
-----------------------
Difference = 13,334.31-12,250 = $1084.31
----------------
Cheers,
Stan H.